# Physical-Layer Machine Learning with Multimode Interferometric Photon Counting

### v1

@@ Abstract
The learning of the physical world relies on sensing and data post-processing. When the signals are weak, multi-dimensional and correlated, the performance of learning is often bottlenecked by the quality of sensors, calling for integrating quantum sensing into the learning of such physical-layer data. An example of such a learning scenario is the stochastic quadrature displacements of electromagnetic fields, modeling optomechanical force sensing, radiofrequency photonic sensing, microwave cavity weak signal sensing and other applications. We propose a unified protocol that combines machine learning with interferometric photon counting to reduce noise and reveal correlations. By applying variational quantum learning with multimode programmable quantum measurements, we enhance signal extraction. Our results show that multimode interferometric photon counting outperforms conventional homodyne detection proposed in prior works for tasks like principal component analysis (PCA) and cross-correlation analysis (CCA), even below vacuum noise levels. To further enhance the performance, We also integrate entanglement-enhanced modules, in the form of squeezed state distribution and anti-squeezing at detection, into the protocol. Combining the multimode interferometric photon counting and multipartite entanglement, the proposed protocol provides a powerful toolbox for learning weak signals.
@@ I Introduction
In contrast, when the physical signals are “weak”, the quality of the data fundamentally limits the performance of the entire learning process, regardless of sophisticated classical or quantum post-processing algorithms. In this regime, quantum sensing [4, 5, 6, 7] becomes central to enhancing the data acquisition precision, speed or bandwidth, through the use of quantum resources such as squeezing, entanglement, or nonclassicality. However, improving data acquisition can be resource-intensive for large data sets, especially when only low-dimensional features of the dataset are ultimately needed. In this regard, integrating the quantum sensing and post-processing into a single module provides a resource-efficient quantum advantage in learning the physical world, as exemplified by the frameworks of Supervised Learning Assisted by an Entangled sensor Network (SLAEN) [8, 9, 10] and quantum computational sensing [11].
@@
A representative example is that of small stochastic quadrature displacements of a resonator (e.g., mechanical oscillator or optical mode) with amplitudes below the vacuum fluctuations. Such signals appear in precision force sensing [13], gravitational wave sensing [14, 15], radiofrequency (RF) detection [12], and dark matter searches [16, 17, 18], and have been increasingly studied in distributed quantum sensing [19]. While many learning strategies in this context adopt squeezing and homodyne detection [8, 9] (and related learning works [20, 21]) — often inspired by coherent displacement sensing protocols — recent results suggest that these tools are strictly suboptimal for weak stochastic signals. In particular, photon counting measurements are known to provide a quantum advantage in the single-mode case [18, 22], but their role in high-dimensional, variationally trained architectures remains largely unexplored.
In this paper, we introduce a framework for physical-layer supervised learning of weak stochastic displacement signals using multimode interferometric photon counting (Fig. 1). Our architecture combines a programmable linear optical (a.k.a., beamsplitter) network and photon counting detectors which are optimized via variational training to solve various learning tasks. As examples, we consider the tasks of PCA and Cross-Correlation Analysis (CCA), which are crucial for data compression and correlation estimation of high-dimensional data.
@@ II.1 Physical-layer quantum data
We assume that the quantum data takes the form of random momentum displacements, denoted by the vector $\vec{\lambda}$, within an $N$-mode quantum system. We model the displacements by a zero-mode random displacement quantum channel, $\Phi_{\bm{V}}$, characterized by the covariance matrix $\bm{V}_{ij}=\mathbb{E}[\vec{\lambda}_{i}\vec{\lambda}_{j}]$ (and $\mathbb{E}[\vec{\lambda}]=0$).II.1II.1II.1The signal can follow a general distribution; however, we focus on signals characterized entirely by their covariance matrix. This assumption simplifies both analytical and numerical analysis. Motivated by optomechanical sensing, we assume all displacements occur along $N$ momentum quadratures, though one may extend our framework to include conjugate position quadratures. Given the momentum operator vector $\hat{\vec{P}}$, the channel acts instance-wise as $\Phi_{\bm{V}}:\hat{\vec{P}}\rightarrow\hat{\vec{P}}+\vec{\lambda}$, where $\vec{\lambda}$ is the stochastic displacement with $\mathbb{E}[\lambda_{i}\lambda_{j}]=\bm{V}_{ij}$. Typically, the quantum data are high-dimensional, encoded across many modes ($N\gg 1$). Furthermore, we assume operation in the quantum regime (i.e., low signal-to-noise ratio), where $|\bm{V}|\ll 1$; that is, the typical magnitude of momentum fluctuations is significantly smaller than vacuum fluctuations, which are of order 1.
@@ II.2 Variational quantum processing
To begin, we focus on the “passive” sensing scenarios. That is, we vary the detection (receiver) while considering the probes being in their vacuum state. This is motivated by weak stochastic signal sensing, where vacuum and photon counting provides a significant advantage over quantum sensing schemes based on homodyne or heterodyne detection [18, 22]. In Section V, we extend towards further quantum advantage from multipartite entanglement via distributed squeezing.
@@
Detectors measure the output probes, yielding estimates of the collective displacement data. These estimates inform the updates to the variational circuit (i.e., the parameters of $\bm{B}$) to minimize a loss function, and thus optimize data processing for tasks like PCA and CCA (see below). This iterative feedback thus trains and refines the variational circuit based on sensor-derived data. In conventional machine learning, the loss function is well-defined, and gradient-based optimization is often effective. However, in quantum systems, measurement-induced uncertainty introduces substantial quantum noise into the loss function, especially in the weak signal case of interest. This noise severely hampers gradient-based optimization, as small parameter updates can lead to disproportionately large fluctuations in the estimated gradient. To address this challenge, our protocol incorporates this inherent poisoning effect by employing a gradient-free optimization algorithm.
We will detail such training process in Section III.
@@
The standard for measuring displacements and quadrature fluctuations is homodyne detection. In this case, the output amplitudes of the variational circuit are directly measured. Notably, this implies that we may perform local homodyne measurements on all modes and then combine the measurement results in classical post-processing, viz., constructing bilinear estimators from the measurement data to learn properties of $\bm{V}$. Consequently, the unitary $\hat{U}_{\rm ml}$ need not be physically implemented, as its action can be emulated in post-processing. Nevertheless, we retain $\hat{U}_{\rm ml}$ in all figures to maintain consistency across measurement protocols.II.2II.2II.2The physical implementation of $\hat{U}_{\rm ml}$ remains relevant in scenarios where, for instance, the number of available homodyne detectors is limited or measurement efficiency is a limiting factor.
@@ II.4 Interesting Problems (PCA and CCA)
In this work, we focus on two specific problems. The first is Principal Component Analysis (PCA), where the objective is to identify the principal component $\vec{w}^{\star}$ of the covariance matrix $\bm{V}$ such that $\vec{w}^{\star}=\arg\max_{\vec{w}}(\vec{w}^{T}\bm{V}\vec{w})$. Hence, $\vec{w}^{\star}\bm{V}\vec{w}^{\star}$ represents the maximal variance of $\bm{V}$. The second problem is what we call Collective Cross-correlation Analysis (CCA), in which, given a fixed reference vector $\vec{u}$, we seek the orthonormal vector $\vec{w}^{\star}$ that maximizes the cross-correlation magnitude, i.e., $\vec{w}^{\star}={\rm argmax}_{\vec{w}}(|\vec{u}^{T}\bm{V}\vec{w}|)$. In both problems, the converged quantum circuit set-up provides vector information on component of interest, while the resulting measurement provides additional amplitude information. Formally, we can define the PCA and CCA variables for a fixed variational circuit as
$\vartheta_{\rm PCA}(\vec{w})\coloneqq\vec{w}^{T}\bm{V}\vec{w}\mbox{\quad and\quad}\vartheta_{\rm CCA}(\vec{w})\coloneqq\vec{u}^{T}\bm{V}\vec{w},$ (II.1)
@@ III.1 Principle Component Analyzer (PCA)
$\hat{L}=-\hat{N}_{w},$ (III.1)
where $\hat{N}_{w}$ is the photon number operator for mode $w$. Since the output state in this mode is effectively thermal with mean energy $\mathbb{E}[\lambda_{w}^{2}]$, we have $\expectationvalue*{\hat{L}}=-\mathbb{E}[\lambda_{w}^{2}]=-\vartheta_{\rm PCA}(\vec{w})$, as desired. The noise fluctuations of the loss function are determined by the variance of the photon number,
$\mathrm{Var}(\hat{N}_{w})=\mathbb{E}[\lambda_{w}^{2}](\mathbb{E}[\lambda_{w}^{2}]+1)\approx\mathbb{E}[\lambda_{w}^{2}],$ (III.2)
where the approximation holds in the weak signal regime.
@@
$\mathrm{Var}(\hat{P}_{w}^{2})=2\mathrm{Var}(\hat{P}_{w})^{2}\approx 1/2,$ (III.4)
@@
For weak stochastic signals, homodyne detection suffers from a comparatively high and irreducible noise floor. To make this concrete, we compare variances of the loss functions for both detection strategies [see Eqs. (III.2) and (III.3)] and observe that
$\frac{{\rm Var}(\hat{L})\big{|}_{\rm count}}{{\rm Var}(\hat{L})\big{|}_{\rm hom}}\approx 2\mathbb{E}[\lambda_{w}^{2}]\ll 1.$ (III.5)
@@ III.2 Collective Cross-Correlation Analysis (CCA)
The signal of interest is the cross-correlation between the two projected modes $\vartheta_{\rm CCA}(\vec{w})=\mathbb{E}[\lambda_{w}\lambda_{u}]=\vec{w}^{T}\bm{V}\vec{u}$. To estimate this quantity at the measurement stage, we adopt an interferometric readout strategy: the two output modes $\lambda_{w}$ and $\lambda_{u}$ are interfered on a balanced (50:50) beamsplitter, followed by photon counting at both output ports. The post-processed difference in photon counts yields access to the cross-correlation — a technique we refer to as interferometric photon counting.III.1III.1III.1A similar strategy applies to homodyne detection, though in that case, “interference” can be implemented in classical post-processing.
@@
$\mathrm{Var}(\hat{N}_{+}-\hat{N}_{-})\approx\mathbb{E}[\lambda_{w}^{2}]+\mathbb{E}[\lambda_{u}^{2}].$ (III.7)
@@
$\mathrm{Var}(\hat{P}_{w}\hat{P}_{u})=\langle\hat{P}_{w}^{2}\rangle\langle\hat{P}_{u}^{2}\rangle-\langle\hat{P}_{w}\hat{P}_{u}\rangle^{2}\approx 1/4,$ (III.9)
@@ III.3 Training methods
To concretely illustrate our methods, we consider the case of maximally correlated noise with the covariance matrix $\bm{V}=M\sigma_{c}^{2}\vec{v}\,\vec{v}^{T}$, where $\vec{v}=(1,\dots,1)/\sqrt{M}$ is a normalized vector and $\sigma_{c}$ characterizes the fluctuation amplitude per mode. At each measurement step, we sample a displacement vector of the form $\vec{\lambda}=\lambda\vec{v}$, with $\lambda\sim\mathcal{N}(0,\sqrt{M}\sigma_{c})$. A single-shot measurement of the circuit output then yields $L_{i}$.
Due to significant fluctuations in the loss observable, we employ particle swarm optimization (PSO), a gradient-free method known for its robustness to noise. The optimization process involves $N_{\rm p}$ parallel particles (or trajectories), each evolving over time via ‘position’ $\vec{w}^{(j)}(t)$ and ‘momentum’ $\vec{d}^{(j)}(t)$. At each epoch $t$, the update rules for particle $j$ are:
@@
$\vec{d}^{(j)}(t+1)$ $=$ $m_{a}\vec{d}^{(j)}(t)+r_{1}\left(\vec{w}^{(j)}(t)-\vec{w}_{\rm best}(t)\right)$ (III.10)
@@
where $m_{a}$ is the inertia coefficient, and $r_{1},r_{2}\in(0,0.5)$ are random generated numbers for each update. The term $\vec{w}_{\rm best}(t)=\arg\min_{w\in\{w^{(j)}(t)|1\leq j\leq N_{\rm p}\}}L[w]$ refers to the best configuration (i.e., with the lowest sampled loss function) found by all particle at time $t$. Similarly, $\vec{w}_{\rm gbest}(t)$ is the best global configuration across all particles and all previous epochs. If $L[\vec{w}_{\rm gbest}(t)]>L[\vec{w}_{\rm best}(t-1)]$, then it is updated with $\vec{w}_{\rm gbest}(t)=\vec{w}_{\rm best}(t)$.
@@
$\vec{w}_{\rm gbest}(t)=(1-g)\,\vec{w}_{\rm gbest}(t-1)+g\,\vec{w}_{\rm best}^{(j)}(t).$ (III.11)
This modification tempers abrupt changes in the optimization landscape and helps stabilize convergence.
Finally, initial positions $\vec{w}^{(j)}(0)$ are chosen randomly, and all initial displacements $\vec{d}^{(j)}(0)$ are set to zero. The first update is set to be $\vec{w}_{\rm gbest}(0)=\vec{w}_{\rm best}(0)$.
Figure 2: Training history of the accuracy of (a) PCA with photon counting, (b) PCA with homodyne detection, (c) CCA with photon counting, and (d) CCA with homodyne detection. Other parameters are $M=21$ and $\sigma_{c}=0.02$. The orange curve shows the accuracy of the global best set up across all time up to $t$, while the blue curve shows the optimal set up at time step $t$.

### v2

@@ Abstract
The learning of the physical world relies on sensing and data post-processing. When the signals are weak, multidimensional and correlated, the performance of learning is often bottlenecked by the quality of sensors, calling for integrating quantum sensing into the learning of such physical-layer data. An example of such a learning scenario is the stochastic quadrature displacements of electromagnetic fields, modeling optomechanical force sensing, radiofrequency photonic sensing, microwave cavity weak signal sensing, and other applications. We propose a unified protocol that combines machine learning with interferometric photon counting to reduce noise and reveal correlations. By applying variational quantum learning with multimode programmable quantum measurements, we enhance signal extraction. Our results show that multimode interferometric photon counting outperforms conventional homodyne detection proposed in prior works for tasks like principal component analysis (PCA) and cross-correlation analysis (CCA), even below vacuum noise levels. To further enhance the performance, we also integrate entanglement-enhanced modules, in the form of squeezed state distribution and anti-squeezing at detection, into the protocol. Combining multimode interferometric photon counting and multipartite entanglement, the proposed protocol provides a powerful toolbox for learning weak signals.
@@ I Introduction
In contrast, when the physical signals are “weak”, the quality of the data fundamentally limits the performance of the entire learning process, regardless of sophisticated classical or quantum post-processing algorithms. In this regime, quantum sensing [4, 5, 6, 7] becomes central to enhancing the data acquisition precision, speed or bandwidth, through the use of quantum resources such as squeezing, entanglement, or nonclassicality. However, improving data acquisition can be resource-intensive for large data sets, especially when only low-dimensional features of the dataset are ultimately needed. In this regard, integrating the quantum sensing and post-processing into a single module provides a resource-efficient quantum advantage in learning the physical world, as exemplified by the frameworks of Supervised Learning Assisted by an Entangled sensor Network (SLAEN) [8, 9, 10] and quantum computational sensing [11, 12, 13].
@@
A representative example is that of small stochastic quadrature displacements of a resonator (e.g., mechanical oscillator or optical mode) with amplitudes below the vacuum fluctuations. Such signals appear in precision force sensing [15], gravitational wave sensing [16, 17], radiofrequency (RF) detection [14], and dark matter searches [18, 19, 20], and have been increasingly studied in distributed quantum sensing [21]. While many learning strategies in this context adopt squeezing and homodyne detection [8, 9] (and related learning works [22, 23]) — often inspired by coherent displacement sensing protocols — recent results suggest that these tools are strictly suboptimal for weak stochastic signals. This motivates the framework of quantum computational sensing, which leverages sequences of quantum sensing, quantum operations, measurements, and optimization to achieve both reduced hardware overhead and lower error probability [13, 12]. In particular, photon counting measurements are known to provide a quantum advantage in the single-mode case [20, 24], but their role in high-dimensional, variationally trained architectures remains largely unexplored.
In this paper, we introduce a framework for physical-layer supervised learning of weak stochastic displacement signals using multimode interferometric photon counting (Fig. 1). Our architecture combines a programmable linear optical (a.k.a., beamsplitter) network and photon counting detectors which are optimized via variational training to solve various learning tasks. Our module thus serves a variational quantum measurement toolbox [25]. As examples, we consider the tasks of PCA and Cross-Correlation Analysis (CCA), which are relevant to data compression and correlation estimation of high-dimensional data [26]. These machine-learning techniques are widely used in experiments, for instance, in image recognition tasks implemented on a nuclear magnetic resonance (NMR) quantum processor [27, 28]. Our framework applies to some of these application scenarios in force and field sensing with mechanical sensors as well as thermal-field imaging, as we discuss further in Section II.4.1.
@@ II.1 Physical-layer quantum data
We assume that the quantum data takes the form of random momentum displacements, denoted by the vector $\vec{\lambda}$, within an $M$-mode quantum system. We model the displacements by a zero-mode random displacement quantum channel, $\Phi_{\bm{V}}$, characterized by the covariance matrix $\bm{V}_{ij}=\mathbb{E}[\vec{\lambda}_{i}\vec{\lambda}_{j}]$ (and $\mathbb{E}[\vec{\lambda}]=0$). [^1] The signal can follow a general distribution; however, we focus on signals characterized entirely by their covariance matrix. This assumption simplifies both analytical and numerical analysis. Motivated by optomechanical sensing, we assume all displacements occur along $M$ momentum quadratures, though one may extend our framework to include conjugate position quadratures. The momentum operator vector is given by $\hat{\vec{P}}\equiv i\left(\hat{\vec{a}}^{\dagger}-\hat{\vec{a}}\right)/\sqrt{2}$, where $\hat{\vec{a}}$ is the photon annihilation operator for the $M$ channels. The channel acts instance-wise as $\Phi_{\bm{V}}:\hat{\vec{P}}\rightarrow\hat{\vec{P}}+\vec{\lambda}$, where $\vec{\lambda}$ is the stochastic displacement with $\mathbb{E}[\lambda_{i}\lambda_{j}]=\bm{V}_{ij}$. Typically, the quantum data are high-dimensional, encoded across many modes ($M\gg 1$). Furthermore, we assume operation in the quantum regime (i.e., low signal-to-noise ratio), where $|\bm{V}|\ll 1$; that is, the typical magnitude of momentum fluctuations is significantly smaller than vacuum fluctuations, which are of order 1.
This framework can be naturally extended to handle simple isotropic displacements affecting both quadratures equally, i.e., identical shifts in $Q$ and $P$, which occurs in RF-photonics [14, 30], optical imaging [31, 32, 33, 34, 35], and microwave cavity settings [18, 19, 36]. In this case, the analysis proceeds analogously to the single-quadrature scenario, with the relevant covariance matrix reflecting the combined displacements. Asymmetric displacements which differ between $Q$ and $P$ are more challenging to address, as photon counting is insensitive to this asymmetry and, thus, separating quadrature-specific fluctuations might require adaptive measurements or additional probe resources.
@@ II.2 Variational quantum processing
To begin, we focus on the “passive” sensing scenarios. That is, we vary the detection (receiver) while considering the probes being in their vacuum state. This is motivated by weak stochastic signal sensing, where vacuum and photon counting provides a significant advantage over quantum sensing schemes based on homodyne or heterodyne detection [20, 24]. In Section VI, we extend towards further quantum advantage from multipartite entanglement via distributed squeezing.
@@
Detectors measure the output probes, yielding estimates of the collective displacement data. These estimates inform the updates to the variational circuit (i.e., the parameters of $\bm{B}$) to minimize a loss function, and thus optimize data processing for tasks like PCA and CCA (see below). This iterative feedback thus trains and refines the variational circuit based on sensor-derived data.
In conventional machine learning, the loss function is well-defined, and gradient-based optimization is often effective. However, in quantum systems, measurement-induced uncertainty introduces substantial quantum noise into the loss function, especially in the weak signal case of interest. This noise severely hampers gradient-based optimization, as small parameter updates can lead to disproportionately large fluctuations in the estimated gradient. To address this challenge, our protocol incorporates this inherent poisoning effect by employing a gradient-free optimization algorithm.
We detail the training process in Section III.
@@
The standard for measuring displacements and quadrature fluctuations is homodyne detection. In this case, the output amplitudes of the variational circuit are directly measured. Notably, this implies that we may perform local homodyne measurements on all modes and then combine the measurement results in classical post-processing, viz., constructing bilinear estimators from the measurement data to learn properties of $\bm{V}$. Consequently, the unitary $\hat{U}_{\rm ML}$ need not be physically implemented, as its action can be emulated in post-processing. Nevertheless, we retain $\hat{U}_{\rm ML}$ in all figures to maintain consistency across measurement protocols. [^2] The physical implementation of $\hat{U}_{\rm ML}$ remains relevant in scenarios where, for instance, the number of available homodyne detectors is limited or measurement efficiency is a limiting factor.
@@
We focus on homodyne detection and photon counting in this work because they are experimentally accessible and widely implemented. Importantly, a variational linear-optical circuit followed by photon-counting detection already constitutes a type of joint non-Gaussian measurement, as local photon counting alone cannot achieve the same performance. Other joint non-Gaussian measurements are possible but are generally more complex and challenging to realize.
Moreover, this measurement scheme is particularly relevant for the tasks studied here, as it is known to be near-optimal for estimating stochastic displacements [37, 20] and for imaging weak thermal fields [33]. Heterodyne detection could serve as a useful benchmark as well, but the expected performance is on par with homodyne due to the inherent linearity of the measurement.
@@ II.4 Interesting Problems (PCA and CCA)
In this work, we focus on two specific problems. The first is Principal Component Analysis (PCA) [38], where the objective is to identify the principal component $\vec{w}^{\star}$ of the covariance matrix $\bm{V}$ such that $\vec{w}^{\star}=\arg\max_{\vec{w}}(\vec{w}^{T}\bm{V}\vec{w})$. Hence, $\vec{w}^{\star}\bm{V}\vec{w}^{\star}$ represents the maximal variance of $\bm{V}$. The second problem is what we call Collective Cross-correlation Analysis (CCA), in which, given a fixed reference vector $\vec{u}$, we seek the orthonormal vector $\vec{w}^{\star}$ that maximizes the cross-correlation magnitude, i.e., $\vec{w}^{\star}={\rm argmax}_{\vec{w}}(|\vec{u}^{T}\bm{V}\vec{w}|)$. In both problems, the converged quantum circuit set-up provides vector information on component of interest, while the resulting measurement provides additional amplitude information. Formally, we can define the PCA and CCA variables for a fixed variational circuit as
$\vartheta_{\rm PCA}(\vec{w})\coloneqq\vec{w}^{T}\bm{V}\vec{w}\mbox{\quad and\quad}\vartheta_{\rm CCA}(\vec{w})\coloneqq\vec{u}^{T}\bm{V}\vec{w},$ (1)
@@
II.4.1 Physical Relevance
Our motivation for focusing on PCA and CCA is that, in many quantum sensing settings, data processing is inseparable from the physical measurement itself: The most informative physical quantities of interest (the signals) often appear in collective or correlated modes, rather than in the raw measurement basis. From this perspective, PCA and CCA are not merely pre-processing tools but form part of the measurement strategy, identifying which physical modes to access and how to combine them. Below we point out representative scenarios where our analysis may be useful.
Mechanical field/force sensing.
In arrays of trapped ions [39] or optomechanics [14, 40, 15], the raw measurement data naturally lives in the position basis of individual oscillators. However, the physically relevant signals, originating from background fluctuating fields or forces, often manifest most strongly in a small number of collective modes, such as the center-of-mass or antisymmetric modes. PCA provides a natural way to extract these dominant modes, while our CCA formulation illustrates how correlations between distinct collective modes can be probed within the same variational framework.
Thermal-field imaging.
Detection scenarios in which the collected field is stochastic and obeys thermal photon statistics arise in many physical settings. Examples include incoherent fluorescence imaging [31, 32, 33, 34, 35], sensing with a microwave cavity network [19], and radio-frequency imaging [14, 30]. In these multimode settings, the raw basis rarely provides direct access to the most informative physical quantities of interest, particularly when the signals are small. As known in the quantum imaging literature [32], the signals may be concentrated in only a few structured or collective modes (e.g., Hermite-Gaussian). Identifying this basis is therefore crucial, and PCA serves as a natural tool for this task.
@@ III.1 Principle Component Analyzer (PCA)
$\hat{L}=-2\hat{N}_{w},$ (2)
where $\hat{N}_{w}$ is the photon number operator for mode $w$. The output state in this mode is an anisotropic thermal state with mean energy $\mathbb{E}[\lambda_{w}^{2}]/2$ — thus, $\expectationvalue*{\hat{L}}=-\mathbb{E}[\lambda_{w}^{2}]=-\vartheta_{\rm PCA}(\vec{w})$, as desired. The noise fluctuations of the loss function are determined by the variance of the photon number,
$\mathrm{Var}(\hat{L})=4\mathrm{Var}(\hat{N}_{w})\approx 2\mathbb{E}[\lambda_{w}^{2}],$ (3)
which holds in the weak-signal regime.
@@
$\mathrm{Var}(\hat{P}_{w}^{2})=2\mathrm{Var}(\hat{P}_{w})^{2}\approx 1/2,$ (5)
@@
For weak stochastic signals, homodyne detection suffers from a comparatively high and irreducible noise floor. To make this concrete, we compare variances of the loss functions for both detection strategies [see Eqs. (3) and (4)] and observe that
$\frac{{\rm Var}(\hat{L})\big|_{\rm count}}{{\rm Var}(\hat{L})\big|_{\rm hom}}\approx 4\mathbb{E}[\lambda_{w}^{2}]\ll 1.$ (6)
@@ III.2 Collective Cross-Correlation Analysis (CCA)
The signal of interest is the cross-correlation between the two projected modes $\vartheta_{\rm CCA}(\vec{w})=\mathbb{E}[\lambda_{w}\lambda_{u}]=\vec{w}^{T}\bm{V}\vec{u}$. To estimate this quantity at the measurement stage, we adopt an interferometric readout strategy: the two output modes $\lambda_{w}$ and $\lambda_{u}$ are interfered on a balanced (50:50) beamsplitter, followed by photon counting at both output ports. The post-processed difference in photon counts yields access to the cross-correlation. [^3] A similar strategy applies to homodyne detection, though in that case, “interference” can be implemented in classical post-processing.
@@
$\mathrm{Var}(\hat{N}_{+}-\hat{N}_{-})\approx\mathbb{E}[\lambda_{w}^{2}]+\mathbb{E}[\lambda_{u}^{2}].$ (8)
@@
$\mathrm{Var}(\hat{P}_{w}\hat{P}_{u})=\langle\hat{P}_{w}^{2}\rangle\langle\hat{P}_{u}^{2}\rangle-\langle\hat{P}_{w}\hat{P}_{u}\rangle^{2}\approx 1/4,$ (10)
@@ III.3 Training methods
To concretely illustrate our methods, we consider the case of maximally correlated noise with the covariance matrix $\bm{V}=2M\sigma_{c}^{2}\vec{v}\,\vec{v}^{T}$, where $\vec{v}=(1,\dots,1)/\sqrt{M}$ is a normalized vector and $\sigma_{c}$ characterizes the fluctuation amplitude per mode. At each measurement step, we sample a displacement vector of the form $\vec{\lambda}=\lambda\vec{v}$, with $\lambda\sim\mathcal{N}(0,\sqrt{2M}\sigma_{c})$. A single-shot measurement of the circuit output then yields $L_{i}$.
There is an unavoidable contribution from finite sampling noise, which cannot be completely eliminated and leads to significant fluctuations in the loss observable. In conventional gradient descent, the gradient error scales as the noise strength divided by the step size. Consequently, attempting to improve accuracy by decreasing the step size paradoxically amplifies the error. To address this limitation, several alternative optimization strategies have been developed that are more robust to noise. One example is particle swarm optimization (PSO), a gradient-free method known for its robustness in noisy environments [41, 42, 43]. Other noise-resilient optimization techniques may also serve as viable alternatives.
In PSO, the optimization process involves $N_{\rm p}$ parallel particles (or trajectories), each evolving over time via ‘position’ $\vec{w}^{(j)}(t)$ and ‘momentum’ $\vec{d}^{(j)}(t)$. At each epoch $t$, the update rules for particle $j$ are:
@@
$\vec{d}^{(j)}(t+1)$ $=$ $m_{a}\vec{d}^{(j)}(t)+r_{1}\left(\vec{w}^{(j)}(t)-\vec{w}_{\rm best}(t)\right)$ (11)
@@
where $m_{a}$ is the inertia coefficient, and $r_{1},r_{2}\in(0,0.5)$ are random generated numbers for each update. The term $\vec{w}_{\rm best}(t)=\arg\min_{w\in\{w^{(j)}(t)|1\leq j\leq N_{\rm p}\}}L[w]$ refers to the best configuration (i.e., with the lowest sampled loss function) found by all particle at time $t$. Similarly, $\vec{w}_{\rm gbest}(t)$ is the best global configuration across all particles and all previous epochs. If $L[\vec{w}_{\rm gbest}(t)]>L[\vec{w}_{\rm best}(t)]$, then it is updated with $\vec{w}_{\rm gbest}(t)=\vec{w}_{\rm best}(t)$.
@@
$\vec{w}_{\rm gbest}(t)=(1-g)\,\vec{w}_{\rm gbest}(t-1)+g\,\vec{w}_{\rm best}^{(j)}(t).$ (12)
This modification tempers abrupt changes in the optimization landscape and helps stabilize convergence. In the ideal limit of infinite sampling, this method becomes unnecessary and can be neglected by setting $g=0$.
In classical machine learning, limited training data often leads to overfitting, which degrades the model’s ability to generalize. In the quantum setting, an additional challenge arises: the loss function cannot be exactly evaluated with a finite number of measurements due to inherent quantum randomness. As a result, an incorrect outcome may occasionally appear with a deceptively low loss, while a correct outcome may yield a spuriously high loss. One simple strategy to mitigate this problem is the introduction of “forgetting” in the learning process [44, 45]. By reducing the influence of outliers that disproportionately disrupt optimization, forgetting ensures continuous parameter updates and improves the average performance of the model.
The initial positions $\vec{w}^{(j)}(0)$ are randomly initialized, with all initial displacements $\vec{d}^{(j)}(0)$ set to zero. The first global best position is defined as $\vec{w}_{\rm gbest}(0)=\vec{w}_{\rm best}(0)$. Throughout the optimization process, we take $\vec{w}_{\rm gbest}^{(j)}(t)$ as the final output instead of $\vec{w}_{\rm best}^{(j)}(t)$.
Figure 2: Training history of the accuracy of (a) PCA with photon counting, (b) PCA with homodyne detection, (c) CCA with photon counting, and (d) CCA with homodyne detection. Other parameters are $M=21$ and $\sigma_{c}=0.02$. The blue solid curve indicates the accuracy of the best-performing setup at each time step $t$, while the orange dashed curve represents the highest accuracy achieved by the global best configuration up to time $t$.